OpenStudy (anonymous):

6 - b/3a for a = 10, b = –9

OpenStudy (anonymous):

\[\Large 6-\frac{ -9 }{3(10) }=6+\frac{ 3 }{ 10 }=\frac{ 60 }{ 10 }+\frac{ 3 }{ 10}=\frac{ 63 }{ 10}\]

OpenStudy (anonymous):

Step 1. Plug in the numbers in the designated letters. Step 2. Simplify